### v2

@@ 2.2.3 Marked graphs and the space $\mathcal{G}_{*}[\mathcal{Y}]$
A sequence $(G_{n},c^{n},y^{n})$ converges locally to $(G,c,y)$ in $\mathcal{G}_{*}[\mathcal{Y}]$ if, for any $k\in\mathbb{N}$ and any $\varepsilon>0$, there exists $N_{k,\varepsilon}\in\mathbb{N}$ sufficiently large so that, for all $n\geq N_{k}$, $B_{G_{n}}(o_{n},k)\cong B_{G}(o,k)$, and that there exists an isomorphism $\varphi$ from $B_{G}(o,k)$ to $B_{G_{n}}(o_{n},k)$ such that, for every $v\in B_{G}(o,k)$, $d(y_{v},y_{\varphi(v)}^{n})<\varepsilon$, and $\sum_{u\in N_{G}^{-}(v)}|c_{uv}-c_{\varphi(u)\varphi(v)}^{n}|<\varepsilon$.
@@
  • (3)
    Following the convention in [28], we use vertex marks to represent the initial conditions $(x_{v})_{v\in V}$, the driving noises $(B_{v})_{v\in V}$, and the solution trajectories $(X_{v})_{v\in V}$.
@@ 2.2.4 Compactness in $\mathcal{G}_{*}[\mathcal{Y}]$
We next state a technical lemma that provides a general method for establishing compactness in $\mathcal{G}_{*}[\mathcal{Y}^{\prime}\times\mathcal{Y}]$ by leveraging compactness criteria in the underlying space $\mathcal{Y}$. The proof of Lemma 2.4 is in Section A.
@@ 2.3 Graph convergence in the local weak sense
For a (possibly random) finite marked graph $(G,y)$, define its empirical distribution associated with the marks as
@@
$\lim_{n\to\infty}\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{\mathcal{C}_{v}(G_{n},c^{n},y^{n})}=\mathcal{L}\left(G,c,y\right)\quad\text{in probability in $\mathcal{P}(G_{*}[\mathcal{Y}])$},$ (2.1)
@@
We refer the reader to [35, Chapter 2] for a detailed discussion on the local convergence of random graphs and examples.
2.4 The space $\mathcal{D}\mathrel{\mathop{\ordinarycolon}}=D([-1,\infty))$ and the $M_{1}$-topology
2.4.1 The space $D([a,b])$
2.4.2 The space $\mathcal{D}\mathrel{\mathop{\ordinarycolon}}=D([-1,\infty))$
Lemma 2.7.
2.4.3 Towards compactness in $\mathcal{D}$
The next result provides a sufficient condition for compactness of subsets of $D([-1,\infty))$ consisting of monotone and uniformly bounded paths. This compactness criterion will be useful in proving tightness of empirical processes arising from our particle system. The proof of the theorem is deferred to Section A.
Theorem 2.8.
$\mathcal{M}((M_{m})_{m\geq 1})$ $\mathrel{\mathop{\ordinarycolon}}=\{f\in D([-1,\infty))\mathrel{\mathop{\ordinarycolon}}\,f=0\,\,\text{on }[-1,0),\,\text{$f$ is monotone and }\sup_{[-1,m]}|f|\leq M_{m},\quad\forall m\geq 1\}.$
In particular, for any $M\in(0,\infty)$, the set
has compact closure in the $M_{1}$ topology.
@@ 3 Main Results and Proofs
We now present the main analytical results of the paper, beginning with the existence and construction of physical solutions to the particle system defined in equations (1.1) and (1.2). As discussed in Section 1, minimal fixed points of the map $\Gamma$ are natural candidates for physical solutions.
@@ 3.1 Physical and minimal solutions
Minimal solution exists, and its $D$-component can be obtained by
@@
$x_{v}+B_{v}(t_{0})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\Gamma^{(N)}[\emptyset]_{t_{0}}\}}\leq 0,$
@@
It is clear that the above construction of minimal solution is pathwise, i.e. there exists a deterministic map $\varphi$ such that the minimal solution $(G,X,D)=\varphi(G,c,x,B)$.
@@
Conditioning on $\sigma_{n}$ and using the reflection principle and the density of the Brownian first hitting time (see [33, Theorem 3.7.1]), we compute:
@@
$(\text{change of variable: }t\mapsto x_{n}^{2}t)\quad$ $=\int_{0}^{\infty}\left(1-2\Phi\left(-\frac{1}{2x_{n}\sqrt{t}}\right)\right)^{n}\,\cdot\frac{1}{t\sqrt{2\pi t}}e^{-\frac{1}{2t}}\,\mathrm{d}t,$
@@
By the Borel–Cantelli lemma, we conclude that $\sigma_{n}\to 0$ almost surely. Therefore, $D_{0+}=V=\mathbb{N}$ with probability $1$ for any solution $(X,D)$. Since $x_{n}>0$ for all $n\geq 0$, any physical solution $(X,D)$ must satisfy $D_{0}=\emptyset$ by the definition of the physicality condition. As a result, the right continuity will be violated and hence there is no physical solution to equations (1.1) and (1.2) with the given initial conditions. ∎
@@ 3.2 Fragility, robustness and default cascades
A vertex $v\in V$ is said to be fragile at time $t$ subject to the configuration $(G,c,x,B)$, if
$0\leq x_{v}+B_{v}(t)\leq w_{v},$
where $w_{v}\mathrel{\mathop{\ordinarycolon}}=\sum_{u\in N_{G}^{-}(v)}c_{uv}$ is the total exposure of vertex $v$ from its neighbors. The set of all vertices that are fragile subject to the configuration $(G,c,x,B)$ at time $t$ is denoted by $F_{t}$.
Definition 3.4.
For every $t\geq 0$, define the following events
$\mathrm{C}_{t}\mathrel{\mathop{\ordinarycolon}}=\left\{F_{t}\text{ contains an infinite weakly connected component}\right\},$
$\mathrm{C}_{[t_{1},t_{2}]}\mathrel{\mathop{\ordinarycolon}}=\left\{\bigcup_{s\in[t_{1},t_{2}]}F_{s}\text{ contains an infinite weakly connected component}\right\},$
where connectivity is considered in the undirected graph obtained from $G$ by ignoring edge orientations. The configuration $(G,c,x,B)$ is said to drive a robust system, if
$\mathbb{P}\left[\,\neg\mathrm{C}_{t},\,\,\forall\,t\geq 0\right]=1.$
It is said to drive a $\delta$-robust system, if
$\mathbb{P}\left[\,\neg\mathrm{C}_{[t,t+\delta]},\,\,\forall\,t\geq 0\right]=1.$
In the next lemma, we outline a few sufficient conditions for $(G,c,x,B)$ to drive $\delta$-robust systems.
@@
The configuration $(G,c,x,B)$ drives $\delta$-robust systems for some $\delta>0$, provided that at least one of the following conditions holds:
@@
    (i.i.d. Brownian noises) $G$ is countably infinite, locally finite and
@@
      $(B_{v})_{v\in V}$ are independent standard Brownian motions and $(x_{v})_{v\in V}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\thicksim}}\mu$.
@@
      $p_{\eta}\mathrel{\mathop{\ordinarycolon}}=\sup_{t\geq 0}\mathbb{P}[x_{v}+B_{v}(t)\in(-\eta,w+\eta)]<p_{c},$ (3.1)
@@
      There exist independent standard Brownian motions $(\tilde{B}_{v})_{v\in V}$ and another continuous process $Z$ that is independent of $(\tilde{B}_{v})_{v\in V}$ such that $B_{v}(t)=\tilde{B}_{v}(t)+Z(t)$, and $(x_{v})_{v\in V}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\thicksim}}\mu$.
@@
    $G=\tilde{G},\quad x_{v}=\lambda_{v}\tilde{x}_{v},\quad c_{uv}=\lambda_{v}\tilde{c}_{uv},\quad B_{v}=\lambda_{v}\tilde{B}_{v}.$
@@
$A^{(v)}_{t,2\delta}\mathrel{\mathop{\ordinarycolon}}=\left\{x_{v}+B_{v}(s)\notin[0,w],\ \forall s\in[t,t+2\delta]\right\}\subset\left\{v\notin F_{s},\ \forall s\in[t,t+2\delta]\right\}.$
To ease the notation, we omit the index $v$ when there is no ambiguity. Notice that for $s>t$, $x_{0}+B_{s}=Z_{t}+\tilde{B}_{s-t}$, with $Z_{t}\mathrel{\mathop{\ordinarycolon}}=x_{0}+B_{t}\thicksim\mu*N(0,t)$, $\tilde{B}_{r}\mathrel{\mathop{\ordinarycolon}}=B_{t+r}-B_{r}\thicksim N(0,r)$ independent of $Z_{t}$. Denote $\mu_{t}=\mu*N(0,t)$. We rewrite $A_{t,\delta}$ as
$A_{t,2\delta}=\{\forall r\in[0,2\delta],\,Z_{t}+\tilde{B}_{r}\notin[0,w]\}.$
@@
$q^{+}(z,2\delta)\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}[\forall r\in[0,2\delta],z+\tilde{B}_{r}\notin[0,w]]=\mathbb{P}\Big{[}\inf_{r\in[0,2\delta]}(z+\tilde{B}_{r})>w\Big{]}=2\Phi(\frac{z-w}{\sqrt{2\delta}})-1$
by the reflection principle. Similarly, for $z\leq-\eta$,
$q^{-}(z,2\delta)\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}[\forall r\in[0,2\delta],z+\tilde{B}_{r}\notin[0,w]]=\mathbb{P}\Big{[}\sup_{r\in[0,2\delta]}(z+\tilde{B}_{r})<0\Big{]}=2\Phi(-\frac{z}{\sqrt{2\delta}})-1.$
@@
$\mathbb{P}[A_{t,2\delta}]$ $=\mathbb{E}\left[\mathbb{P}[Z_{t}+\tilde{B}_{r}\notin[0,w],\,\forall r\in[0,2\delta]\,|\,Z_{t}]\right]\geq\mathbb{E}[q(Z_{t},2\delta)]$
$=\int_{-\infty}^{-\eta}q^{-}(z,2\delta)\mu_{t}(z)dz+\int_{w+\eta}^{\infty}q^{+}(z,2\delta)\mu_{t}(z)dz$
$\geq(2\Phi(\frac{\eta}{\sqrt{2\delta}})-1)\left(1-\mathbb{P}[Z_{t}\in(-\eta,w+\eta)]\right).$
From equation (3.1), $p_{\eta}=\sup_{t>0}\mathbb{P}[Z_{t}\in(-\eta,w+\eta)]\leq p_{c}-\varepsilon$. Then for every $t>0$,
$\mathbb{P}[A_{t,2\delta}]\geq(2\Phi(\frac{\eta}{\sqrt{2\delta}})-1)(1-p_{c}+\varepsilon).$
If $\delta$ is chosen such that
$0<\sqrt{2\delta}\leq\frac{\eta}{\Phi^{-1}(\frac{1-p_{c}+\varepsilon/2}{1-p_{c}+\varepsilon})},$
then $\mathbb{P}[A_{t,2\delta}]\geq 1-p_{c}$ for every $t>0$. Since the $(x_{v},B_{v})$ are i.i.d., the events $A^{(v)}_{t,2\delta}$ are independent across $v\in V$. Therefore, by the percolation theory (see [20, Section 1.4-1.6]),
$\mathbb{P}\left[\neg\mathrm{C}_{[t,t+2\delta]}\right]=1,\quad\forall t\geq 0.$
By taking a countable union over $t=0,\delta,2\delta,$..., we obtain
$\mathbb{P}\left[\bigcap_{n=0}^{\infty}\neg\mathrm{C}_{[n\delta,(n+2)\delta]}\right]=1,$
@@
We then turn to show some technical results concerning minimal solutions for which the configuration $(G,c,x,B)$ drives robust systems:
@@
Let the configuration $(G,c,x,B)$ drive robust systems and let $(X,D)$ be the minimal solution. Then
@@
$\sigma_{w}\mathrel{\mathop{\ordinarycolon}}=\inf\Big{\{}s\geq t\mathrel{\mathop{\ordinarycolon}}\,x_{w}+B_{w}(s)-\sum_{u\in N_{G}^{-}(v)}c_{uw}\mathbf{1}_{\{u\in D_{t}\}}\leq 0\Big{\}}>t.$
@@
$\sigma_{u}\mathrel{\mathop{\ordinarycolon}}=\inf\Big{\{}s\geq t\mathrel{\mathop{\ordinarycolon}}\,x_{u}+B_{u}(s)-\sum_{w\in N_{G}^{-}(u)}c_{wu}\leq 0\Big{\}}>t.$
@@
$\inf_{s\in[0,t]}(x_{w}+B_{w}(s))\geq\inf_{s\in[0,t]}X_{w}(s)>0\quad\text{and}$
$x_{w}+B_{w}(s)\geq x_{w}+B_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in D_{t}\}}>0,\quad\forall s\in(t,\sigma),$
@@
$x_{w}+B_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in\Gamma^{(N)}[\emptyset]_{s}\}}\geq x_{w}+B_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in D_{t}\}}>0$
@@
$\inf_{s\in[0,t]}\Big{(}x_{w}+B_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in\Gamma^{(N)}[\emptyset]_{s}\}}\Big{)}\geq\inf_{s\in[0,t]}\Big{(}x_{w}+B_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in D_{s}\}}\Big{)}>0.$
@@
Let the configuration $(G,c,x,B)$ drive robust systems. Then the minimal solution is physical.
@@
$D_{t}^{(N+1)}$ $\mathrel{\mathop{\ordinarycolon}}=D_{t}^{(N)}\cup\Big{\{}v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{(N)}\}}\leq 0\Big{\}}$
$\subset D_{t}^{(N)}\cup\Big{\{}v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}}\leq 0\Big{\}}\subset D_{t}.$
@@
$x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t-}\}}>0,\quad\forall v\notin D_{t-}\quad\text{implies }\quad D_{t}=D_{t-}.$

### v3

@@ 2.1 Elementary notations
We denote by $\mathcal{D}\mathrel{\mathop{\ordinarycolon}}=D([-1,\infty))$ the space of functions $f\mathrel{\mathop{\ordinarycolon}}[-1,\infty)\to\mathbb{R}$ that are right-continuous at all $t\in[-1,\infty)$ and have left limits at all $t\in(-1,\infty)$. The space $\mathcal{D}$ will be endowed with the Skorokhod $M_{1}$ topology throughout the paper. A brief technical introduction to the $M_{1}$ topology can be found in Appendix A.
@@ 2.2.3 Marked graphs and the space $\mathcal{G}_{*}[\mathcal{Y}]$
A sequence $(G_{n},c^{n},y^{n})$ converges locally to $(G,c,y)$ in $\mathcal{G}_{*}[\mathcal{Y}]$ if, for any $k\in\mathbb{N}$ and any $\varepsilon>0$, there exists $N_{k,\varepsilon}\in\mathbb{N}$ sufficiently large so that, for all $n\geq N_{k,\varepsilon}$, $B_{G_{n}}(o_{n},k)\cong B_{G}(o,k)$, and that there exists an isomorphism $\varphi$ from $B_{G}(o,k)$ to $B_{G_{n}}(o_{n},k)$ such that, for every $v\in B_{G}(o,k)$, $d(y_{v},y_{\varphi(v)}^{n})<\varepsilon$, and $\sum_{u\in N_{G}^{-}(v)}|c_{uv}-c_{\varphi(u)\varphi(v)}^{n}|<\varepsilon$.
@@
  3. (c)
    Following the convention in [26], we use vertex marks to represent the initial conditions $(x_{v})_{v\in V}$, the driving noises $(Z_{v})_{v\in V}$, and the solution trajectories $(X_{v})_{v\in V}$.
@@ 2.2.4 Compactness in $\mathcal{G}_{*}[\mathcal{Y}]$
We next state a technical lemma that provides a general method for establishing compactness in $\mathcal{G}_{*}[\mathcal{Y}^{\prime}\times\mathcal{Y}]$ by leveraging compactness criteria in the underlying space $\mathcal{Y}$. The proof of Lemma 2.4 is in Section B.
@@ 2.3 Graph convergence in the local weak sense
For a (possibly random) finite marked graph $(G,y)$, the empirical distribution associated with the marks is defined as
@@
$\lim_{n\to\infty}\frac{1}{|G_{n}|}\sum_{v\in G_{n}}\delta_{\mathcal{C}_{v}(G_{n},c^{n},y^{n})}=\mathcal{L}\left(G,c,y\right)\quad\text{in probability in $\mathcal{P}(\mathcal{G}_{*}[\mathcal{Y}])$},$ (2.1)
@@
We refer the reader to [33, Chapter 2] for a detailed discussion on the local convergence of random graphs and examples.
@@ 3 Main Results and Proofs
We now present the main analytical results of the paper, beginning with the existence and construction of physical solutions to the particle system defined in equations (1.1). As discussed in Section 1, minimal fixed points of the map $\Gamma$ are natural candidates for physical solutions.
@@ 3.1 Physical and minimal solutions
The minimal solution exists, and its $D$-component can be obtained by
@@
$x_{v}+Z_{v}(t_{0})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in\Gamma^{(N)}[\emptyset]_{t_{0}}\}}\leq 0,$
@@
It is clear that the above construction of minimal solution is pathwise, i.e. there exists a deterministic map $\varphi$ such that the minimal solution $(G,X,D)=\varphi(G,c,x,Z)$.
@@
Conditioning on $\sigma_{n}$ and using the reflection principle and the density of the Brownian first hitting time (see [31, Theorem 3.7.1]), we compute:
@@
$(\text{change of variable: }t\mapsto x_{n}^{2}t)$ $=\int_{0}^{\infty}\left(1-2\Phi\left(-\frac{1}{2x_{n}\sqrt{t}}\right)\right)^{n}\,\cdot\frac{1}{t\sqrt{2\pi t}}e^{-\frac{1}{2t}}\,\mathrm{d}t,$
@@
By the Borel–Cantelli lemma, we conclude that $\sigma_{n}\to 0$ almost surely. Therefore, $D_{0+}=V=\mathbb{N}$ with probability $1$ for any solution $(X,D)$. Since $x_{n}>0$ for all $n\geq 0$, any physical solution $(X,D)$ must satisfy $D_{0}=\emptyset$ by the definition of the physicality condition. As a result, the right continuity will be violated and hence there is no physical solution to equations (1.1) with the given initial conditions. ∎
@@ 3.2 Fragility, robustness and default cascades
Recall the definition of fragility and robustness mentioned in Section 1.
@@
A vertex $v\in V$ is said to be fragile at time $t$ subject to the configuration $(G,c,x,Z)$, if
$v\in F_{t}\mathrel{\mathop{\ordinarycolon}}=\Big\{v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+Z_{v}(t-)\geq 0\quad\text{and}\quad x_{v}+Z_{v}(t)\leq\sum_{u\in N_{G}^{-}(v)}c_{uv}\Big\}.$
Remark 3.4.
Let $(X,D)$ be a solution to the system (1.1). Then it is straightforward to see that $v\in F_{\tau_{v}}$ provided that $\tau_{v}<\infty$, where
$\tau_{v}\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}\,X_{v}(t)\leq 0\}=\inf\{t\geq 0\mathrel{\mathop{\ordinarycolon}}\,v\in D_{t}\}.$
In the next lemma, we outline a few sufficient conditions for $(G,c,x,Z)$ to drive $\delta$-robust systems.
@@
The configuration $(G,c,x,Z)$ drives $\delta$-robust systems for some $\delta>0$, provided that at least one of the following conditions holds:
@@
    (i.i.d. Lévy noises) $G$ is countably infinite, locally finite and
@@
      $(Z_{v})_{v\in V}$ are i.i.d. Lévy processes and $(x_{v})_{v\in V}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\thicksim}}\mu$.
@@
      $p_{\eta}\mathrel{\mathop{\ordinarycolon}}=\sup_{t\geq 0}\mathbb{P}[x_{v}+Z_{v}(t-)\in(-\eta,w+\eta)]<p_{c},$ (3.1)
@@
      There exist i.i.d. Lévy processes $(\tilde{Z}_{v})_{v\in V}$ and another càdlàg process $Z^{0}$ that is independent of $(\tilde{Z}_{v})_{v\in V}$ such that $Z_{v}(t)=\tilde{Z}_{v}(t)+Z^{0}(t)$, and $(x_{v})_{v\in V}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\thicksim}}\mu$.
@@
    $G=\tilde{G},\quad x_{v}=\lambda_{v}\tilde{x}_{v},\quad c_{uv}=\lambda_{v}\tilde{c}_{uv},\quad Z_{v}=\lambda_{v}\tilde{Z}_{v}.$
@@
$A^{(v)}_{t,2\delta}\mathrel{\mathop{\ordinarycolon}}=\left\{x_{v}+Z_{v}(s-)<0\,\text{ or }\,x_{v}+Z_{v}(s)>w,\ \forall s\in[t,t+2\delta]\right\}\subset\left\{v\notin F_{s},\ \forall s\in[t,t+2\delta]\right\}.$
To ease the notation, we omit the index $v$ when there is no ambiguity. Notice that for $s>t$, $x+Z(s)=(x+Z(t-))+\tilde{Z}(s-t)$ where $\tilde{Z}(r)\mathrel{\mathop{\ordinarycolon}}=Z(t+r)-Z(t-)$ is another Lévy process with the same distribution as $Z$, and is independent of x+Z(t)x+Z(t-). Denote by $\mu_{t-}$ the law of x+Z(t)x+Z(t-). We rewrite $A_{t,2\delta}$ as
$A_{t,2\delta}=\{\forall r\in[0,2\delta],\,x+Z(t-)+\tilde{Z}(r-)<0\,\text{ or }x+Z(t-)+\tilde{Z}(r)>w\}.$
@@
$q^{+}(z,2\delta)\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}[\forall r\in[0,2\delta],z+\tilde{Z}(r)>w]\geq\mathbb{P}\Big[\inf_{r\in[0,2\delta]}\tilde{Z}(r)>-\eta\Big].$
Similarly, for $z\leq-\eta$,
$q^{-}(z,2\delta)\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}[\forall r\in[0,2\delta],z+\tilde{Z}(r)<0]\geq\mathbb{P}\Big[\sup_{r\in[0,2\delta]}\tilde{Z}(r)<\eta\Big].$
@@
$\mathbb{P}[A_{t,2\delta}]$ $=\mathbb{E}\left[\mathbb{P}[A_{t,2\delta}|\,x+Z(t-)]\right]\geq\mathbb{E}[q(x+Z(t-),2\delta)]$
$=\int_{-\infty}^{-\eta}q^{-}(z,2\delta)\mu_{t-}(dz)+\int_{w+\eta}^{\infty}q^{+}(z,2\delta)\mu_{t-}(dz)$
$\geq\mathbb{P}\Big[\sup_{r\in[0,2\delta]}|\tilde{Z}(r)|<\eta\Big]\left(1-\mathbb{P}[x+Z(t-)\in(-\eta,w+\eta)]\right).$
From equation (3.1), $p_{\eta}=\sup_{t>0}\mathbb{P}[x+Z(t-)\in(-\eta,w+\eta)]\leq p_{c}-\varepsilon$. In addition, as $\tilde{Z}$ is a Lévy process, we can choose $\delta$ sufficiently small such that
$\mathbb{P}\Big[\sup_{r\in[0,2\delta]}|\tilde{Z}(r)|<\eta\Big]\geq\frac{1-p_{c}}{1-p_{c}+\varepsilon}.$
Then $\mathbb{P}[A_{t,2\delta}]\geq 1-p_{c}$ for every $t>0$. Since the $(x_{v},Z_{v})$ are i.i.d., the events $A^{(v)}_{t,2\delta}$ are independent across $v\in V$. Therefore, by the percolation theory (see [20, Section 1.4-1.6]),
$\mathbb{P}\Bigg[\bigcup_{s\in[t,t+2\delta]}F_{s}\text{ contains no infinite weakly connected component}\Bigg]=1,\quad\forall t\geq 0.$
By taking a countable intersection over $t=0,\delta,2\delta,$..., we obtain
$\mathbb{P}\Bigg[\bigcap_{n=0}^{\infty}\Big\{\bigcup_{s\in[n\delta,(n+2)\delta]}F_{s}\text{ contains no infinite weakly connected component}\Big\}\Bigg]=1,$
@@
We then turn to show some technical results concerning minimal solutions for which the configuration $(G,c,x,Z)$ drives robust systems:
@@
Let the configuration $(G,c,x,Z)$ drive robust systems and let $(X,D)$ be the minimal solution. Then
@@
$\sigma_{w}\mathrel{\mathop{\ordinarycolon}}=\inf\Big\{s\geq t\mathrel{\mathop{\ordinarycolon}}\,x_{w}+Z_{w}(s)-\sum_{u\in N_{G}^{-}(v)}c_{uw}\mathbf{1}_{\{u\in D_{t}\}}\leq 0\Big\}>t.$
@@
$\sigma_{u}\mathrel{\mathop{\ordinarycolon}}=\inf\Big\{s\geq t\mathrel{\mathop{\ordinarycolon}}\,x_{u}+Z_{u}(s)-\sum_{w\in N_{G}^{-}(u)}c_{wu}\leq 0\Big\}>t.$
@@
$\inf_{s\in[0,t]}(x_{w}+Z_{w}(s))\geq\inf_{s\in[0,t]}X_{w}(s)>0\quad\text{and}$
$x_{w}+Z_{w}(s)\geq x_{w}+Z_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in D_{t}\}}>0,\quad\forall s\in(t,\sigma),$
@@
$x_{w}+Z_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in\Gamma^{(N)}[\emptyset]_{s}\}}\geq x_{w}+Z_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in D_{t}\}}>0$
@@
$\inf_{s\in[0,t]}\Big(x_{w}+Z_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in\Gamma^{(N)}[\emptyset]_{s}\}}\Big)$
$\geq\inf_{s\in[0,t]}\Big(x_{w}+Z_{w}(s)-\sum_{u\in N_{G}^{-}(w)}c_{uw}\mathbf{1}_{\{u\in D_{s}\}}\Big)>0.$
@@
Let the configuration $(G,c,x,Z)$ drive robust systems. Then the minimal solution is physical.
@@
$D_{t}^{(N+1)}$ $\mathrel{\mathop{\ordinarycolon}}=D_{t}^{(N)}\cup\Big\{v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{(N)}\}}\leq 0\Big\}$
$\subset D_{t}^{(N)}\cup\Big\{v\in V\mathrel{\mathop{\ordinarycolon}}\,x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}}\leq 0\Big\}\subset D_{t}.$
@@
$x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t-}\}}>0,\quad\forall v\notin D_{t-}\quad\text{implies }\quad D_{t}=D_{t-}.$